Reports 1-1 of 1 Clear search Modify search
Infrastructures (Air Conditioning)
soldani - 16:24 Wednesday 14 June 2023 (60550) Print this report
INJ AHU frequency test
This morning, upon request and in agreement with M.Tringali and A.Chiummo, a test was carried out for the INJ AHU unit, setting various frequencies of the inverters of the air supply and return motors;
list of actions (in local time):
- 12,02: supply inverter from 22 to 17 Hz (in manual mode)/return inverter 21 Hz (in manual mode)
- 12,12: supply inverter 17 Hz (in manual mode)/return inverter 17 Hz (in manual mode)
- 12,25: both inverter (and AHU) off
- 12,31: AHU on with old set: supply inverter 25 Hz (in manual mode)/return inverter in automatic mode
Comments to this report:
Tringali, Chiummo - 19:10 Wednesday 14 June 2023 (60553) Print this report

The purpose of this test was to continue the investigation (#59807) of the bumps at ~16.9Hz and  ~18.8Hz related to LB and EIB, respectively. 

Figure 1  shows the variation of the supply (*FREQ_OUT*) and return (*FREQ_IN*) fan frequency, the temperature and overpressure in the laser room during the test. We did not observe significant temperature variation inside the laser laboratory.
In Figure 2, we report the spectra of the microphone and accelerometers on the laser bench (top row) and on the external injection bench (bottom row). The slowdown of the AHU did not produce any effect on the amplitude of the two bumps.
However, we observe a further excitation of  a resonance at  ~13 Hz visible in all acoustic and seismic sensors when the supply frequency is shifted to 22 Hz, see Figure 3.
This increased resonance is also visible in the quadrants of the beam pointing control mostly on the vertical dof, Figure 4.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×